OpenStudy (anonymous):

How arctan 3/4 is equal to 36degre and 52 min

OpenStudy (anonymous):

Convert the arctan of 3/4 radians to degrees. \[\frac{180 \tan ^{-1}\left(\frac{3}{4}\right)}{\pi }=36.8699{}^{\circ} \]Convert 0.8699 degrees to minutes.\[.8699(60 \min )=52.194 \min \]